OpenStudy (anonymous):

HELP PLEASE? WILL MEDAL! (: Generate the first five terms of this sequence: f(1) = 1 and f(2) = 2, f(n) = f(n-1) + f (n-2) for n>2 A) 1, 2, 3, 4, 5 B) 1, 2, 3, 5, 7 C) 1, 2, 3, 5, 8 D) 1, 3, 5, 7, 9

OpenStudy (anonymous):

f(3)=f(2)+f(1)=2+1=3 you can find other terms like this.

OpenStudy (anonymous):

Okay so, f(1) = f(2) and f(n-1) + f(n-2)?

OpenStudy (anonymous):

I don't even know if i'm making any sense rn, but i'm gonna try to understand haha!

Nnesha (nnesha):

n= term that you hve to find so to find 3rd term you should replace n with 3

Nnesha (nnesha):

\[\huge\rm f(\color{Red}{n}-1) = f(\color{ReD}{3}-1)\] solve the parentheses 3-1 =2 that's how he got f(2)

Nnesha (nnesha):

\[\huge\rm f(\color{Red}{3}) = f(\color{ReD}{3}-1)+f(\color{Red}{3}-2)\] solve the parentheses 3-1 =2 that's how he got f(2) 3-1 is one so f(3) = f(2)+f(1) now replace f(2) and f(1) with their values

Nnesha (nnesha):

given f(1)= 1 f(2) = 2 so substitute f(1) for 1 f(2) for 2

OpenStudy (anonymous):

So where do those 1, and 2 go? I don't see f(2) or f(1) in the equation that you put above which is the f(3) = f(3-1) + f(3-2)

OpenStudy (anonymous):

@Nnesha

Nnesha (nnesha):

yes replace f(1) and f(2) given is : f(1) = 1 f(2) =2 so plug them in \[\huge\rm f(\color{Red}{3}) = f(\color{ReD}{3}-1)+f(\color{Red}{3}-2)\] f(1) = 1 f(2) =2 so plug them in \[\huge\rm f(\color{Red}{3}) = \color{ReD}1+\color{Red}{2}\] 1+2 = 3 so f(3) = 3 now replace n with 4 to find 4th term